Yard Designs (YD) experienced the following events in Year 1, its first year of operation:

  1. On October 1, Year 1, YD collected $40,800 for consulting services it agreed to provide during the coming year.
  2. Adjusted the accounts to reflect the amount of consulting service revenue recognized in Year 1.

  
Required
Based on this information alone:
a. Record the events under an accounting equation.
b. Prepare an income statement, balance sheet, and statement of cash flows for the Year 1 accounting period.
c. Ignoring all other future events, what is the amount of service revenue that would be recognized in Year 2?

Record the events under an accounting equation. (Do not round intermediate calculations. Enter any decreases to account balances with a minus sign.) YARD DESIGNS Accounting Equations for the Year 1 Assets Liabilities + Stockholders Equity Retained Earnings Event CashUnearned Revenue |=| Event Adj

Prepare an income statement for the Year 1 accounting period. (Do not round intermediate calculations.) YARD DESIGNS Income Statement For the Year Ended December 31, Year 1 0

Prepare an income statement for the Year 1 accounting period. (Do not round intermediate calculations.) YARD DESIGNS Income Statement For the Year Ended December 31, Year 1 0

YARD DESIGNS Balance Sheet As of December 31, Year 1 Assets Total assets Liabilities Total liabilities Stockholders Equity Total stockholders equity Total liabilities and stockholders equity

YARD DESIGNS Statement of Cash Flows For the Year Ended December 31, Year 1 Cash flows from operating activities Net cash flow from operating activities Cash flows from investing activities Cash flows from financing activities Net change in cash Ending cash balance

ignoring all other future events, what is the amount of service revenue that would be recognized in Year 2? (Do not round intermediate calculations.) Service revenue to be recognized in Year 2 < Req B Stmt of Cash Flows Req C
Answer #1
a.
Event Assets = Liabilities + S/E
Cash = Unearned revenue + Retained earnings
event 40800 = 40800 +
Adj = -10200 + 10200
40800 = 30600 + 10200
40800*3/12=10200
b)
Yard Designs Income Statement
For the Year Ended December 31, 2018
Revenue 10200
Expense 0
Net Income 10200
c)
Yard Designs Balance Sheet
As of December 31, 2018
Assets
Cash 40800
Total Assets 40800
Liabilities
Unearned Revenue 30600
Total Liabilities 30600
Stockholders’ Equity
Retained Earnings 10200
Total Stockholders’ Equity 10200
Total Liab. and Stockholders’ Equity 40800
d)
Yard Designs Statement of Cash Flows
For the Year Ended December 31, 2018
Cash Flows From Operating Activities:
Cash Receipt from Revenue 40800
Net Cash Flow from Operating Activities 40800
Cash Flows From Investing Activities -
Cash Flows From Financing Activities:
Net Change in Cash 40800
Plus: Beginning Cash Balance 0
Ending Cash Balance 40800
Ques c
Nine months of unearned revenue from 2018 will be recognized in 2019: $40,500 x 9/12 = $30,600